OpenStudy (anonymous):

this one is complicated... x^2+2x-14=10. completing the square.

OpenStudy (anonymous):

and solve.

OpenStudy (anonymous):

(x+6)(x-2)=0 x=-6,2

OpenStudy (anonymous):

add 25: then x^2+2x+1=(x+1)^2

OpenStudy (anonymous):

yah you are right add 25

OpenStudy (anonymous):

x= -6 and 4?

OpenStudy (anonymous):

s x=-6,4 sorry for typing mistake